9. A small manufacturer that makes clothespins and other household products buys new injection molding equipment for a cost of $500,000. This will allow the manufacturer to make more clothespins in the same amount of time with an estimated increase in sales of 25%. If the manufacturer currently makes 75 tons of clothespins per year, which sell at $18,000 per ton, what will be the increase in revenue next year from the new equipment?

A) $337,500

B) $303,750

C) $125,000

D) $837,500
Answer #1

A) $337,500

Current sale quantity = 75 tons

Percentage increase in sale quantity = 25%

New sale quantity = 75*(1.25) = 93.75 tons

Increase in sale quantity = 93.75 - 75 = 18.75 tons

Rate per ton = $18,000

Thus,

Increase in revenue next year would be:

= 18.000 * 18.75

= $337.500

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.